two days at bush gardens and 3 days at universal studios cost $420 while 4 days at bush gardens and 2 days at universal studios cost $472. what was the cost per day for universal studios
Answer by ikleyn(52834)

.

    2B + 3U = 420,    (1)

    4B + 2U = 472.    (2)


Multiply equation (1) by 2 (both sides). Keep equation (2) as is.


    4B + 6U = 840,    (1')

    4B + 2U = 472.    (2')


Now subtract equation (2') from equation (1').


You will get


         4U = 840 - 472 = 368

          U             = 368/4 = 92  dollars.    ANSWER

Solved.